3. Given utility function U= X0.5Y0.5 where PX = 12 Birr, Birr, PY = 4 Birr and the income of the consumer is, M= 240 Birr.
a. Find the utility maximizing combinations of X and Y.

b. Calculate marginal rate of substitution of X for Y (MRSX,Y) at equilibrium and interpret your result.
1
Expert's answer
2020-05-19T09:53:11-0400

a)


"MU_x=\frac{y^{0.5}}{2x^{0.5}}"


"MU_y=\frac{x^{0.5}}{2y^{0.5}}"


"12x+4y=240"


"\frac{MU_x}{p_x}=\frac {MU_y}{p_y}"


"\frac{y^{0.5}}{24x^{0.5}}=\frac{x^{0.5}}{8y^{0.5}}"


"y=3x"


"x=10, y=30"

b)


"MRS_{x,y}=\frac{MU_x}{MU_y}"


"MRS_{x,y}=\frac{y}{x}=\frac{30}{10}=3"